# Learning Hierarchical Sparse Transform Coding of 3DGS

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) supports fast, high quality, novel view synthesis but has a heavy memory footprint, making the compression of its model crucial. Current state-of-the-art (SOTA) 3DGS compression methods adopt an anchor-based architecture that pairs the Scaffold-GS representation with conditional entropy coding. However, these methods forego the analysis-synthesis transform, a vital mechanism in visual data compression. As a result, redundancy remains intact in the signal and its removal is left to the entropy coder, which computationally overburdens the entropy coding module, increasing coding latency. Even with added complexity thorough redundancy removal is a task unsuited to an entropy coder. To fix this critical omission, we introduce a Sparsity-guided Hierarchical Transform Coding (SHTC) method, the first study on the end-to-end learned neural transform coding of 3DGS. SHTC applies KLT to decorrelate intra-anchor attributes, followed by quantization and entropy coding, and then compresses KLT residuals with a low-complexity, scene-adaptive neural transform. Aided by the sparsity prior and deep unfolding technique, the learned transform uses only a few trainable parameters, reducing the memory usage. Overall, SHTC achieves an appreciably improved R–D performance and at the same time higher decoding speed over SOTA. Its prior-guided, parameter-efficient design may also inspire low-complexity neural image and video codecs. Our code will be released at here.
Figure 1: Illustration of Scaffold-GS and anchor-based compression methods.
@@ 1 Introduction
In recent years, 3D Gaussian Splatting (3DGS) [kerbl20233d] has advanced computer vision and graphics by enabling real-time, high-quality, novel-view synthesis. The cost is substantial storage and bandwidth overheads incurred by storing a very large number of Gaussian primitives, making 3DGS compression an important research topic. Current state-of-the-art (SOTA) methods follow an anchor-based paradigm that pairs Scaffold-GS [lu2024scaffold] with context-based entropy coding to exploit spatial correlations and improve compression ratios [chen2024hac, wang2024contextgs, liu2024hemgs, zhan2025catdgs, chen2025hac++]. However, these methods forego end-to-end optimized analysis-synthesis transforms that remove redundancy before entropy coding. Without such an analysis transform, high-dimensional correlations persist and the inherent sparsity is left underutilized, forcing complex causal entropy models to compensate for the omission. Yet even large entropy models struggle to accurately capture inter-token dependencies, leading to suboptimal rate–distortion (R–D) performance. Additionally, because the 3DGS compression system operates in the “compress once, decompress many times” mode, decoding speed is critical; using complex entropy models inevitably increases decoding latency and degrades the client-side user experience.
To address aforementioned omission and drawbacks, we integrate an end-to-end optimized analysis–synthesis transform into the 3DGS compression pipeline, thereby removing the most of redundancy before entropy coding. Despite the promise of this paradigm and its maturity in neural image and video compression [balle2018variational, minnen2018joint, cheng2020learned, li2024neural], applying it to 3DGS compression requires overcoming two domain-specific challenges. First, point clouds in 3DGS are unstructured: unlike pixels on a grid, neighboring Gaussians (or anchors) cannot be indexed directly. While K-nearest neighbors (KNN) provides a way to define local neighborhoods on such irregular data, the spatial locations of Gaussians (or anchors) are variables in the R–D optimization and change during training, making per-iteration recomputation of KNN neighborhoods computationally prohibitive. Second, because each scene uses a dedicated transform whose parameters must be transmitted in the bitstream, the parameter budget must remain small; achieving strong performance under this constraint requires careful architectural design.
To tackle these challenges, we propose sparsity-guided hierarchical transform coding (SHTC), which (i) avoids spatial KNN by restricting transforms to channel-wise operations for computational efficiency and (ii) uses a two-layer architecture for parameter efficiency. Layer 1 applies the Karhunen-Loève Transform (KLT) for decorrelation and energy compaction, yielding channels with uneven energy distribution. Transmitting only the principal coefficients causes information loss and limits the achievable rendering quality, whereas coding all coefficients is rate-inefficient. To strike a balance between these extremes, we introduce a neural refinement module in Layer 2, which serves as a refinement layer. It compensates for truncation-induced information loss by transmitting the KLT residual with modest additional rate, using a more flexible strategy than direct coefficient coding. Because the residual components are predominantly small (with many values near zero), we model them as sparse. For such sparse signals, compressed sensing theory shows that the full signal can be reliably recovered from a small number of linear measurements rather than coding all coefficients directly [donoho2006compressed]. Motivated by this, we design the refinement layer following compressed-sensing principles. At the encoder, the residuals are projected into a low-dimensional latent space via a learnable linear transform that acts as a measurement matrix, and at the decoder they are reconstructed by unfolding the classical Iterative Shrinkage–Thresholding Algorithm (ISTA) [zibulevsky2010l1] into a sequence of network layers [gregor2010learning, zhang2020deep, zhang2018ista]. This sparse-recovery-inspired design enables accurate residual reconstruction with low parameter cost.
In conclusion, our key contribution is the introduction of the first end-to-end optimized transform coding framework for 3DGS compression, delivering substantial R-D gains with minimal additional parameters and runtime overhead. Across three standard datasets (Mip-NeRF 360 [barron2022mip], Tanks&Temples [knapitsch2017tanks], and DeepBlending [hedman2018deep]), our method reduces the memory footprint by $123.5\times$, $216.7\times$, and $705.3\times$ relative to vanilla 3DGS, respectively, while matching its rendering quality. In addition, our parameter-efficient design strategy, which combines a sparsity prior with deep unfolding, may offer a potential blueprint for developing low-complexity neural image and video codecs.
@@
2 Related Work
Following the taxonomy of [ali2025compression], we group existing 3DGS compression methods into (i) unstructured approaches, which employ model compression techniques such as pruning [ali2024elmgs, ali2024trimming, fan2024lightgaussian, niemeyer2024radsplat, lee2024compact, hanson2025pup] and quantization [girish2024eagles, navaneet2024compgs, wang2024end] for compression, and (ii) structured approaches [morgenstern2024compact, lee2025compression, lu2024scaffold], which exploit the spatial relationship among Gaussian primitives for compression. Because anchor-based designs are especially prevalent within structured approaches and our method builds on this architecture, we dedicate this section to anchor-based methods. Due to space limitations, a more detailed literature review is provided in the supplementary material.
Among anchor-based 3DGS compression methods, Scaffold-GS [lu2024scaffold] is the representative baseline. Scaffold-GS introduces a hierarchical structure during training: a sparse set of anchors serve as reference points from which a dense set of Gaussians are generated. As shown on the left side of Fig. 1, each anchor is parameterized by the position $\mathbf{p}$, latent feature $\mathbf{f}$, scaling factor $\mathbf{l}$, and $K$ learnable offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$. The scaling factor $\mathbf{l}$ together with the offsets determine the spatial distribution of the generated Gaussians. For each generated Gaussian, the view-dependent color, opacity, and covariance (parameterized by quaternion + scale) are predicted from the anchor’s latent feature through lightweight MLPs.
Although Scaffold-GS is more compact than vanilla 3DGS, it is essentially a 3DGS compaction scheme: anchors and their attributes are still stored as raw 32-bit floats without quantization or entropy coding, which inherently limits the achievable compression ratio. SOGS reduces the dimensionality of anchor features; however, this remains 3DGS compaction rather than true compression, so the gains are still limited [zhang2025sogs]. To address these limitations, recent work introduces context-based conditional entropy models that reduce statistical redundancy across anchor attributes, achieving significantly higher compression ratios. Among these methods, HAC was introduced early and serves as a foundational baseline. As illustrated in the right panel of Fig. 1, HAC uses multi-scale hash grids as a hyperprior [balle2018variational]. Given an anchor’s position, it queries these hash grids to obtain contextual features, which an MLP maps to the quantization step size $q_{s}$, the Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$. HAC++ [chen2025hac++] enhances this design by (i) splitting each anchor feature vector into non-overlapping subgroups and encoding them sequentially with a channel-wise autoregressive model (ChARM) [minnen2020channel], and (ii) replacing a single Gaussian with a Gaussian mixture model to model more complex densities. CAT-3DGS [zhan2025catdgs] also adopts ChARM but substitutes hash-grid hyperpriors with triplanes. Both ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs] follow a spatially sequential scheme: they use previously decoded anchors as conditioning context in a spatial autoregressive model (SARM) to encode the next anchors. These approaches represent the current state-of-the-art compression performance for 3DGS compression.
3 Research Motivation and Challenges
In this section, we analyze the importance and challenges of implementing end-to-end optimized transform coding in 3DGS compression.
@@ 3.1 Motivation
Anchor-based compression methods, built on Scaffold-GS, achieve high compression ratios by quantizing and entropy-coding anchor attributes. Let $\mathbf{x}$ denote an anchor-attribute vector. With entropy coding, the ideal rate equals the source entropy $H_{q}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log q(\mathbf{x})]$, but the ground truth distribution $q(\mathbf{x})$ is unknown. In practice, one can approximate the unknown distribution $q(\cdot)$ with a parametric model $p(\cdot)$ (often modeled as Gaussian), so the actual rate is the cross-entropy
$R=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log p(\mathbf{x})]=H_{q}+D_{KL}(q||p).$ (1)
where $D_{KL}(q||p)$ measures the extra bits caused by the mismatch between $q(\cdot)$ and $p(\cdot)$. Since anchor attributes are correlated, anchor-based methods use autoregressive models to exploit these correlations and reduce $D_{KL}$, thereby tightening the cross-entropy rate. For intra-vector correlation, $\mathbf{x}$ can be partitioned into groups of size $g$ and coded sequentially with the probability model $p(\mathbf{x}_{i:i+g}|\mathbf{x}_{<i})$ (e.g., ChARM as used in HAC++ [chen2025hac++] and CAT-3DGS [zhan2025catdgs]); for spatial correlation, anchors are coded causally with the probability model $p(\mathbf{x}^{(i)}|\{\mathbf{x}^{(j)}\}_{j\in\mathcal{C}_{i}})$ where $\mathcal{C}_{i}$ denotes the causal neighbors of anchor-$i$ (SARM as in ContextGS [wang2024contextgs] and HEMGS [liu2024hemgs]). While more complex entropy models can improve compression by exploiting data correlations during entropy coding, they also increase decoding latency. In 3DGS’s “compress once, decompress many times” pattern, decoding latency on the client side is critical; complex entropy models slow down decoding and degrade user experience. Moreover, even with more complex entropy models, fully capturing these dependencies remains difficult, leading to sub-optimal compression performance. These limitations arise from the absence of end-to-end optimized analysis–synthesis transforms. This motivates introducing such transforms into the compression pipeline: by decorrelating the signal before entropy coding, redundancy is reduced and much simpler, more efficient entropy models can be used. As shown in the first row of Fig. 2, HAC and HAC++ anchor features exhibit strong inter-channel correlations. Applying a channel-wise KLT largely removes these correlations, yielding decorrelated coefficients and thereby paving the way for the use of simpler entropy models.
@@
Applying an analysis transform before entropy coding decorrelates the data and compacts its energy, which benefits lossy compression, while a paired synthesis transform reconstructs the anchor attributes for rendering. When the transforms and entropy models are jointly optimized, this transform coding paradigm typically achieves superior R–D performance, as demonstrated in neural image compression [balle2016end, balle2018variational, minnen2018joint, cheng2020learned, he2022elic, zou2022devil, zhu2021transformer, liu2023learned, ali2024towards]. Motivated by these successes, extending this paradigm to 3DGS compression is a natural choice.
3.2 Challenges
Despite the promise of end-to-end optimized transform coding, to the best of our knowledge, no existing work has applied such transforms to 3DGS compression. Extending end-to-end optimized transform coding from neural image compression to 3DGS is, however, non-trivial due to two domain-specific challenges.
@@ 3.2 Challenges
The first challenge stems from the irregular, unordered nature of 3DGS anchor points, which lack a natural neighborhood structure. Although neighborhoods can be defined via KNN, repeated neighbor searches in dense point clouds are computationally expensive. This issue is exacerbated by the fact that anchor positions are optimization variables that evolve during training, so the KNN graph must be rebuilt throughout optimization rather than constructed once, significantly increasing the overall training cost.
The second challenge arises from a fundamental difference between 3DGS compression and neural image/video codecs. In standard neural codecs, the analysis–synthesis transforms are trained once, shared across inputs, and omitted from the bitstream; consequently, large-capacity transforms can be used to pursue better R-D trade-offs. By contrast, 3DGS relies on scene-specific transforms whose parameters must be transmitted, imposing a strict parameter budget. Under this constraint, strong R–D performance hinges on architectural efficiency rather than sheer model size. Historically, early neural image codecs with relatively simple transforms [balle2016end, balle2018variational] failed to surpass traditional codecs, whereas later models with substantially more complex transforms did [cheng2020learned]. This evolution highlights the difficulty of designing neural transforms that are both parameter-efficient and capable of strong R-D performance.
These two challenges call for careful architectural design to balance R-D gains against additional computational and parameter costs. To address them, we propose SHTC, a transform-coding framework for 3DGS that is both computationally and parameter efficient, achieving strong R-D performance under a tight parameter budget. Technical details of SHTC are provided in Sec. 4.
4 SHTC: Sparsity-guided Hierarchical Transform Coding
[…]
4.1 Layer 1: Base Layer

### v3

# Learning Hierarchical Sparse Transform Coding for 3DGS Compression

@@ Abstract
Current 3DGS compression methods largely forego the neural analysis-synthesis transform, which is a crucial component in learned signal compression systems. As a result, redundancy removal is left solely to the entropy coder, overburdening the entropy coding module and reducing rate-distortion (R-D) performance. To fix this critical omission, we propose a training-time transform coding (TTC) method that adds the analysis-synthesis transform and optimizes it jointly with the 3DGS representation and entropy model. Concretely, we adopt a hierarchical design: a channel-wise KLT for decorrelation and energy compaction, followed by a sparsity-aware neural transform that reconstructs the KLT residuals with minimal parameter and computational overhead. Experiments show that our method delivers strong R–D performance with fast decoding, offering a favorable BD-rate–decoding-time trade-off over SOTA 3DGS compressors.
Machine Learning, ICML
Figure 1: Overview of 3DGS compression paradigms. Methods are categorized into four paradigms: (1) pruning and quantization, (2) anchor-based structure with adaptive entropy coding, (3) post-training transform coding (PTC) on a frozen 3DGS, and (4) training-time transform coding (TTC), introduced by us, which is the first to jointly optimize 3DGS, context model and the transform for improved R-D performance. Within TTC, we further propose SHTC, a lightweight transform architecture that delivers strong R-D performance with minimal parameter and computation overhead.
@@ 1 Introduction
3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) enables real-time, high-quality novel-view synthesis, but its many Gaussian primitives create significant storage and bandwidth overheads, motivating 3DGS compression. As summarized in Fig. 1, existing 3DGS compression methods can be broadly grouped into unstructured approaches (e.g., pruning and quantization) and structured, entropy-coded pipelines. Within the structured category, most competitive compressors either adopt the anchor-based paradigm by pairing Scaffold-GS (Lu et al., 2024) with context-based entropy coding, or perform post-training transform coding (PTC) on a fixed 3DGS. In anchor-based methods, end-to-end analysis–synthesis transform coding is absent, so redundancy is not sufficiently removed before entropy coding. This leaves high-dimensional dependencies and sparsity underexploited, shifting the burden to increasingly complex entropy models (e.g., HAC++ (Chen et al., 2025c), ContextGS (Wang et al., 2024b)), which can still yield suboptimal rate-distortion (R–D) performance and higher decoding latency. Although a few prior works recognize the importance of transforms (Morgenstern et al., 2024; Lee et al., 2025b), they apply them only post-hoc to a fixed 3DGS, decoupling transform from 3DGS representation learning and preventing mutual adaptation under a unified R–D objective, which often limits compression gains. To address these limitations, we advocate a new paradigm, training-time transform coding (TTC), in which analysis–synthesis transforms are learned during training, allowing the 3DGS representation, entropy model, and transforms to be jointly optimized under a unified R-D objective.
Unlike neural image/video compression where the transform is shared across images and incurs no per-image overhead, TTC uses scene-specific transforms that must be transmitted as part of the bitstream. From the minimum description length (MDL) perspective (Rissanen, 1978), TTC should ideally balance the transform cost and the coding cost of the transformed 3DGS attributes, i.e., $L(M)+L(D|M)$, where $L(M)$ denotes the number of bits to describe the scene-specific transform and $L(D|M)$ is the code length for encoding the transformed 3DGS attributes. In principle, increasing transform complexity can reduce $L(D|M)$, but simultaneously increases $L(M)$, potentially canceling out the compression gain.
Optimizing the full MDL objective for TTC is highly challenging, because it requires non-trivial discrete–continuous optimization over the transform’s parameters and architecture. MDL optimization problem is only well posed for a given the model class. As the attainable trade-off between $L(M)$ and $L(D|M)$ is with respect to the chosen transform family, search-based solvers such as NAS (Elsken et al., 2019) can only explore what this search space permits. Thus, a more important and unanswered question is how to design an appropriate transform architecture for scene-specific 3DGS compression. Once such an architecture defines a suitable transform family as the search space, future work can readily apply existing NAS to search for MDL-efficient designs. In this work, we focus on this prerequisite step and propose an MDL-aware transform architecture that is parameter- and computation-efficient.
To instantiate this MDL-aware, parameter- and computation-efficient transform design within TTC, we propose sparsity-guided hierarchical transform coding (SHTC). SHTC avoids computationally prohibitive repeated spatial KNN graph construction by restricting transforms to the channel domain, and uses a lightweight two-layer hierarchy to keep parameter and computational overhead low. In Layer 1, we apply the Karhunen-Loève Transform (KLT) to decorrelate channels and compact energy, yielding coefficients with highly uneven energy distribution. Coding only the principal coefficients reduces rate but introduces truncation error, whereas coding all coefficients is rate-inefficient. To balance this trade-off, Layer 2 serves as a refinement layer and uses a neural transform to compress the KLT residual, compensating for truncation-induced information loss with modest additional rate. Since the residual is typically compressible with many near-zero entries, we inject this sparsity prior as an inductive bias into the refinement layer, enabling effective residual coding with very few parameters. Concretely, we adopt a compressed-sensing-inspired refinement that encodes the residual with a small set of learned linear measurements and reconstructs it using a lightweight deep-unfolding decoder (Gregor and LeCun, 2010; Zhang et al., 2020; Zhang and Ghanem, 2018).
Overall, we introduce TTC, which enables joint optimization of the 3DGS representation, the entropy model, and the transforms. Within this paradigm, we propose SHTC, a parameter- and computation-efficient transform architecture that achieves substantial R–D gains with minimal decoding-time overhead and moderate training-time cost; moreover, Fig. 2 shows that it is empirically Pareto-optimal in the BD-rate–decoding-time trade-off compared with current state-of-the-art 3DGS compressors. In addition, our parameter-efficient design strategy, which combines a sparsity prior with deep unfolding, may offer a potential blueprint for developing low-complexity neural image and video codecs.
Figure 2: Trade-off between decoding time and BD-rate (lower is better), with marker size indicating training time for HAC, ContextGS, CAT-3DGS, HAC++, and Ours.
Figure 3: Illustration of (a) Scaffold-GS and (b) an anchor-based compression pipeline, using HAC as an example.
2 Preliminary and Related Work
In this work, to achieve the best compression performance, SHTC is integrated into anchor-based frameworks to enable the joint optimization of 3DGS, the entropy model, and the transforms. Since these anchor-based frameworks are built upon Scaffold-GS (Lu et al., 2024), we first recap the Scaffold-GS representation and then review how recent anchor-based methods further compress it. Due to space constraints, a more comprehensive introduction for the unstructured compression methods is deferred to Appendix G.1; a dedicated review of post-training transform coding in existing 3DGS compression approaches is presented in Appendix G.2.
Scaffold-GS introduces a hierarchical structure during training: a sparse set of anchors serve as reference points from which a dense set of Gaussians are generated. As shown on the left side of Fig. 3, each anchor is parameterized by the position $\mathbf{p}$, latent feature $\mathbf{f}$, scaling factor $\mathbf{l}$, and $K$ learnable offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$. The scaling factor $\mathbf{l}$ together with the offsets $\{\mathbf{O}_{i}\}_{i=1}^{K}$ determine the spatial distribution of the generated Gaussians. For each generated Gaussian, the view-dependent color, opacity, and covariance (parameterized by quaternion + scale) are predicted from the anchor’s latent feature through lightweight MLPs.
Although Scaffold-GS is more compact than vanilla 3DGS, it primarily performs representation compaction. Anchor parameters and attributes are still stored as raw 32-bit floats, without quantization or entropy coding, which fundamentally limits the achievable bitrate reduction. SOGS reduces the dimensionality of anchor features (Zhang et al., 2025), but it still falls into the category of compaction rather than learned compression, and the gains remain limited. To overcome these limitations, recent anchor-based methods introduce context-based conditional entropy models to quantize and entropy-code anchor attributes, substantially improving compression ratios. HAC is an early representative baseline in this line of work. As illustrated in the right panel of Fig. 3, HAC employs multi-scale hash grids as a hyperprior (Ballé et al., 2018). Given an anchor position, it queries the hash grids to obtain contextual features, which are then mapped by an MLP to the quantization step size $q_{s}$ and the Gaussian likelihood parameters $(\boldsymbol{\mu},\boldsymbol{\sigma})$. Building on HAC, subsequent methods improve the rate–distortion performance by adopting more expressive entropy models, such as channel-wise autoregressive model (ChARM) (Minnen and Singh, 2020) in HAC++ (Chen et al., 2025c) and CAT-3DGS (Zhan et al., 2025), as well as spatial autoregressive model (SARM) in ContextGS (Wang et al., 2024b) and HEMGS (Liu et al., 2024a). These methods represent the current state-of-the-art compression performance for 3DGS compression.
@@ 2 Preliminary and Related Work
3 Method
This section outlines the motivation and challenges for transitioning from anchor-based compression and PTC to our proposed TTC paradigm. We then present our main technical contribution, SHTC.
@@ 3.1 Motivation
Let $\mathbf{x}$ denote the anchor-attribute vector to be compressed. Under ideal entropy coding, the minimum achievable rate equals the source entropy $H_{q}=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log q(\mathbf{x})]$, but the true distribution $q(\mathbf{x})$ is typically unknown. In practice, one can approximate the unknown distribution $q(\cdot)$ with a parametric model $p(\cdot)$ (often modeled as Gaussian), leading to the cross-entropy rate
$R=\mathbb{E}_{\mathbf{x}\sim q(\mathbf{x})}[-\log p(\mathbf{x})]=H_{q}+D_{KL}(q||p)$ (1)
where $D_{KL}(q||p)$ measures the extra bits due to model mismatch. Anchor attributes exhibit strong statistical dependencies; assuming independence across dimensions/channels therefore increases mismatch and inflates the rate. Existing anchor-based methods mitigate this issue by using correlation-aware entropy models (e.g., ChARM or SARM) to better capture dependencies and reduce $D_{KL}$. However, complex entropy models typically incur higher decoding cost (often due to sequential/context-dependent decoding), which is undesirable in 3DGS’s “compress once, decompress many times” setting where client-side decoding latency is critical. Moreover, fully modeling complex dependencies remains challenging even with sophisticated entropy models, limiting compression gains. A complementary approach is to introduce an analysis transform and decorrelate $\mathbf{x}$ before entropy coding. By reducing redundancy in the representation, we can rely on simpler and faster entropy models without sacrificing rate performance. As illustrated in the first row of Fig. 4, HAC and HAC++ features show pronounced inter-channel correlations; a more complex entropy model (as in HAC++) models these dependencies but does not remove them from the signal. In contrast, applying a channel-wise KLT substantially suppresses inter-channel correlations, producing near-decorrelated coefficients and paving the way for the use of simpler entropy models.
@@
PTC vs. TTC: decoupled vs. co-adaptive optimization.
Applying an analysis transform before entropy coding can decorrelate the signal and concentrate its energy into fewer coefficients, which improves the R-D efficiency of lossy compression. A paired synthesis transform then reconstructs the anchor attributes for rendering. A few prior works have also explored analysis-ynthesis transforms (Morgenstern et al., 2024; Lee et al., 2025b), but they typically follow the PTC paradigm (see Fig. 1): applying a fixed, pre-defined transform codec to a frozen 3DGS model in the post-training stage (details in Sec. G.2). This two-stage, decoupled design prevents mutual adaptation between the transform and the underlying 3DGS representation, often leading to sub-optimal R-D performance. In contrast, we advocate the TTC paradigm, where the transforms and entropy model are trained jointly with the 3DGS representation, allowing them to co-adapt and yielding a more compressible representation without sacrificing rendering quality.
3.2 Challenges and Design Considerations
TTC is an open-ended paradigm with many candidate transforms (e.g., DCT/DWT, generic MLPs, and graph-based spatial transforms). A natural question then arises: what kinds of transforms are appropriate? To answer this question, we need to understand two domain-specific challenges in 3DGS compression.
@@ 3.2 Challenges and Design Considerations
The first challenge stems from the irregular, unordered nature of anchor/Gaussian primitives, which lack a natural neighborhood structure. Although neighborhoods can be defined via K-Nearest Neighbors (KNN), neighbor searches in dense point clouds are expensive. Moreover, anchor/Gaussian positions are optimization variables and thus change during training, precluding a one-time KNN construction and requiring repeated graph rebuilding, which substantially increases training cost. This makes graph-based spatial transforms prohibitively expensive to integrate into end-to-end 3DGS training.
The second challenge arises from a fundamental difference between 3DGS compression and neural image/video codecs (Ballé et al., 2018; Minnen et al., 2018; Cheng et al., 2020; He et al., 2022; Li et al., 2024). Standard neural codecs employ a single analysis-synthesis transform shared across inputs, and its parameters are not included in the bitstream; this allows the use of parameter-heavy transforms to improve the R–D performance. In contrast, under the TTC paradigm, 3DGS relies on scene-specific transforms whose parameters must be transmitted, imposing a strict parameter budget. Consequently, R-D performance hinges on architectural efficiency and cannot be improved simply by increasing the number of model parameters. Under a tight parameter budget, purely linear transforms often lack sufficient expressive power, while purely MLP-based transforms may be ineffective without suitable inductive bias.
In light of these constraints, we propose SHTC. SHTC circumvents the first challenge by restricting the transform to be channel-wise, i.e., without spatial interactions, making it computation-friendly. To remain effective under a low parameter budget, SHTC adopts a hierarchical architecture, with interpretable components. Details are provided in Sec. 3.3.
3.3 SHTC: Sparsity-guided Hierarchical Transform Coding